OpenStudy (anonymous):

what two whole numbers is the square root of 35 between?

OpenStudy (e.mccormick):

Two ways you could do this: take the square root in a calculator, see what that is close to OR start doing squares and see when you pass 35 because the one just before is one and the one just after is the other.

OpenStudy (e.mccormick):

\(2^2=4\) Is 4 more than 35? No. Then 2 is not the whole number above \(\sqrt{35}\). \(3^2=9\) Is 9 more than 35? No. Then 3 is not the whole number above \(\sqrt{35}\). \(4^2=16\) Is 16 more than 35? No. Then 4 is not the whole number above \(\sqrt{35}\). You keep doing that until you get something above 35. Then you know what is right above, and the previous one is right below.

OpenStudy (ybarrap):

x^2 < 35 < y^2 what's the maximum whole number, x^2, that is less than 35 what's the minimum whole number, y^2, that is greater than 35 then x and y are the numbers you need.